# Pushing the Limits of the Reactive Affine Shaker Algorithm to Higher Dimensions

### v1

@@ 3.1 Implementation of the anisotropic search box
The search region $\mathcal{R}$ is implemented as a box defined by $d$ independent base vectors ($\boldsymbol{b}_{1}\ldots\boldsymbol{b}_{d}$), where $d$ is the number of dimensions of the search domain. The search region $\boldsymbol{\Delta}$ is determined as a random linear combination of the base vector with iid coefficients uniformly distributed in $[-1,1]$:
@@
Shape modifications are implemented as affine transformations of these vectors by contracting (if the step fails) or dilating (if the step succeeds) their components along the direction of $\boldsymbol{\Delta}$ as shown in Fig. 2. Analytically:
$\forall j\quad\boldsymbol{b}_{j}\leftarrow\left(\boldsymbol{I}+(\rho-1)\frac{\boldsymbol{\Delta}\boldsymbol{\Delta}^{T}}{\|\boldsymbol{\Delta}\|^{2}}\right)\boldsymbol{b}_{j},$ (1)
where $\rho$ controls the amount of contraction ($10<\rho<$ 1) or dilation ($\rho>1$) of the base vectors along the direction $\boldsymbol{\Delta}$.
@@ 4.2 Comparison with state-of-the-art heuristics
Mopta08 [6] — a 124-parameter vehicle design problem where the objective is the minimization of the vehicle’s mass; the original problem is subject several constraints transformed into soft penalties.
@@ 4.3.1 Isotropic search region
Figure 7: Effect of the double-shot strategy (top left) compared with a single-shot step (top right) on a high-dimensionality function. For reference, the bottom chart reports the same effect on the low-dimensionality Rosenbrock 2D function.
@@ 0.A.1 Random directions tend to be mutually orthogonal
Figure 8: Average angle between random vectors in $d$ dimensions and expected number of double-shot successes for a random displacement depending on the search box radius $r(B^{\prime})$ relative to the average curvature of the function’s isosurface.
@@
The first row of Table 8 reports the experimental average angle between randomly generated vectors in a $d$-dimensional space (average of $10^{4}$ tests).
For $d=2$ the true average angle is $45\degree$, but for higher and higher dimensionalities the average tends to $90\degree$. In the simplifying hypothesis that the function’s gradient is unknown and that we generate a move in a random direction, this means that in high dimensions there is a progressively smaller probability of moving along the gradient.
@@ 0.A.2 Improving moves tend to be scarce
Consider the situation described in Fig. 9. Let $\hat{\boldsymbol{x}}\in\mathbb{R}^{d}$ be the current point during a local search for the minimum of function $f:D\to\mathbb{R}$ on domain $D\subseteq\mathbb{R}^{d}$.
@@
The isosurface has dimension $d-1$ and is embedded in $\mathbb{R}^{d}$. Its locally divides the domain into two sides, the “good” one where $f(\boldsymbol{x})<f(\hat{\boldsymbol{x}})$ (left side of the isoline in Fig. 9) and the “bad” one where $f(\boldsymbol{x})>f(\hat{\boldsymbol{x}})$. If $f$ is locally convex and smooth enough, we can approximate the curvature radius of the iso-surface at point $\hat{\boldsymbol{x}}$ by considering the radius of the largest sphere $B$ tangent to the iso-surface in $\hat{\boldsymbol{x}}$ and lying on the same side of it [^2] For $d>2$, we actually have different curvature radii on different directions; let us consider the smallest one..

### v2

@@ 3.1 Implementation of the anisotropic search box
The search region $\mathcal{R}$ is implemented as a box defined by $d$ independent base vectors ($\boldsymbol{b}_{1}\ldots\boldsymbol{b}_{d}$), where $d$ is the number of dimensions of the search domain. The search region $\boldsymbol{\Delta}$ is determined as a random linear combination of the base vector with independent coefficients uniformly distributed in $[-1,1]$:
@@
Shape modifications are implemented as affine transformations of these vectors by contracting (if the step fails) or dilating (if the step succeeds) their components along the direction of $\boldsymbol{\Delta}$ as shown in Fig. 2. Analytically, we want to add to every base vector $\boldsymbol{b}_{j}$ a contribution in the direction of $\boldsymbol{\Delta}$ proportional to the base vector’s size:
$(\rho-1)\frac{\boldsymbol{\Delta}\cdot\boldsymbol{b}_{j}}{\|\boldsymbol{\Delta}\|^{2}}\boldsymbol{\Delta},$
where $\rho$ controls the amount of contraction ($10<\rho<$ 1, making the contribution opposite wrt to the projection of $\boldsymbol{b}_{j}$ along $\boldsymbol{\Delta}$) or dilation ($\rho>1$, making the contribution positive along the projection). This can be rewritten as the following linear transformation:
$\forall j\quad\boldsymbol{b}_{j}\leftarrow A\boldsymbol{b}_{j},\qquad\text{where}\qquad A=\boldsymbol{I}+(\rho-1)\frac{\boldsymbol{\Delta}\boldsymbol{\Delta}^{T}}{\|\boldsymbol{\Delta}\|^{2}}.$ (1)
@@ 4.2 Comparison with state-of-the-art heuristics
Mopta08 [6] — a 124-parameter vehicle design problem where the objective is the minimization of the vehicle’s mass; the original problem is subject to several constraints transformed into soft penalties.
@@ 4.3.1 Isotropic search region
Figure 7: Effect of the double-shot strategy (top right) compared with a single-shot step (top left) on a high-dimensionality function. For reference, the bottom chart reports the same effect on the low-dimensionality Rosenbrock 2D function.
@@ 0.A.1 Random directions tend to be mutually orthogonal
Table 1: Average angle between random vectors in $d$ dimensions (top row) and expected number of double-shot successes for a random displacement depending on the search box radius $r(B^{\prime})$ relative to the average curvature of the function’s isosurface.
@@
Following [3], the average angle $\bar{\theta}_{d}$ in radians between two random directions in $d\geq 2$ dimensions is
$\bar{\theta}_{d}=\frac{J_{d-2}}{I_{d-2}},$
where the numerator and denominator are recursively defined as
$I_{d}=\begin{cases}\frac{\pi}{2}&\text{if $d=0$}\\ 1&\text{if $d=1$}\\ \frac{(d-1)I_{d-2}}{d}&\text{if $d>1$,}\end{cases}\qquad J_{d}=\begin{cases}\frac{\pi^{2}}{8}&\text{if $d=0$}\\ 1&\text{if $d=1$}\\ \frac{(d-1)J_{d-2}}{d}+\frac{1}{d^{2}}&\text{if $d>1$.}\end{cases}$
The first row of Table 1 reports the values of $\bar{\theta}$ computed for an increasing number of dimensions.
For $d=2$ the average angle is $45\degree$, but for higher and higher dimensionalities the average tends to $90\degree$. In the simplifying hypothesis that the function’s gradient is unknown and that we generate a move in a random direction, this means that in high dimensions there is a progressively smaller probability of moving along the gradient.
@@ 0.A.2 Improving moves tend to be scarce
The effectivenes of the double-shot strategy in the limit for small search boxes was proved in [3]. However, given a desired success probability, as the number of search-space dimensions grows the required search box size might become impractically small. Consider the situation described in Fig. 8. Let $\hat{\boldsymbol{x}}\in\mathbb{R}^{d}$ be the current point during a local search for the minimum of function $f:D\to\mathbb{R}$ on domain $D\subseteq\mathbb{R}^{d}$.
@@
The isosurface has dimension $d-1$ and is embedded in $\mathbb{R}^{d}$. It locally divides the domain into two sides, the “good” one where $f(\boldsymbol{x})<f(\hat{\boldsymbol{x}})$ (left side of the isoline in Fig. 8) and the “bad” one where $f(\boldsymbol{x})>f(\hat{\boldsymbol{x}})$. If $f$ is locally convex and smooth enough, we can approximate the curvature radius of the iso-surface at point $\hat{\boldsymbol{x}}$ by considering the radius of the largest sphere $B$ tangent to the iso-surface in $\hat{\boldsymbol{x}}$ and lying on the same side of it [^2] For $d>2$, we actually have different curvature radii on different directions; let us consider the smallest one..